11. Suppose you throw a ball vertically downward with a speed of 49 m/s What would be the speed of the ball one second later? 39 m/s. Show calculation: (g = 9.8 m/s) 49 m/s. ) 58.8 m/s. () 10 m/s. () 10 m/s2. 12. What distance will it have travelled after one second? Show calculation: 13. A rifle bullet is fired horizontally at the same instant another bullet is allowed to drop from rest at th Which bullet strikes the earth first? () The bullet from the rifle. () The bullet allowed to drop. () Both strike at the same instant. 14. A man standing on a bridge throws a stone horizontally with a speed of 20 m/s. The stone hits the water below 3 s later. The bridge is () 44.1 m high. () 60.1 m high. () 30 m high. C) 20 m high Show calculation: (g 9.8 m/s At what horizontal distance does the stone hit the ground?
Answer #1

From the given question,

11. Initial speed(u)=49 m/s

Acceleration(a)=9.8 m/s^2

Time(t)=1 sec

Final velocity(v) = u+at

=49+(9.8)(1)

=58.8 m/s

12.Distance travelled in 1 sec(s)= ut +(1/2)(a)(t^2)

=(49)(1)+(1/2)(9.8)(1^2)

=58.8 m

13. As the initial vertical velocity of both bullets is same ( 0 m/s), they both will strike at same instant.

14. Initial vertical velocity(u)= 0 m/s

Acceleration due to gravity(a)=9.8 m/s^2

Time= 3 s

Distance travelled(s)=ut+(1/2)at^2

=(0)(3)+(1/2)(9.8)(3^2)

=44.1m

Horizontal distance stone hits= horizontal velocity*time

=20*3

=60 m